An ion with a charge of +3.2×10−19 C is in a region where a uniform electric field of 5×104 V/m
is perpendicular to a uniform magnetic field of 0.8T. If its acceleration is zero then its speed
must be:
A. 0
B. 1.6 × 104 m/s
C. 4.0 × 104 m/s
D. 6.3 × 104 m/s
E. any value but 0
Answer #1

speed of the ion is v = E /B

                                  = 5 *10^4 m / 0.8 T

                                  = 6.3 *10^4 m/s